OpenStudy (anonymous):

find the trigonometric limits. lim (1/t^2)sin^2(t/2) as t approaches to 0.

OpenStudy (zarkon):

1/4

OpenStudy (anonymous):

give me the procedure how u did

OpenStudy (zarkon):

do you know how to calculate \[\lim_{t\to0}\frac{\sin(t/2)}{t}\] ?

OpenStudy (anonymous):

no.

OpenStudy (zarkon):

can you do this ... \[\lim_{t\to0}\frac{\sin(t)}{t}\]

OpenStudy (anonymous):

isn't it 1?

OpenStudy (zarkon):

yes

OpenStudy (zarkon):

How about this then... \[\lim_{t\to0}\frac{\sin(t/2)}{t}=\lim_{t\to0}\frac{1}{2}\frac{\sin(t/2)}{t/2}\]

OpenStudy (anonymous):

why do you multiply 1/2 in front of sin(t/2)/t/2?

OpenStudy (zarkon):

\[\lim_{t\to0}\frac{\sin(t/2)}{t}=\lim_{t\to0}\frac{\sin(t/2)}{2\cdot t/2}=\lim_{t\to0}\frac{1}{2}\frac{\sin(t/2)}{t/2}\]

OpenStudy (anonymous):

thank you.

OpenStudy (zarkon):

no problem